3x-2y=-18
4x+5y=22

Answer by Boreal(15235)

I assume a solution for x and y is wanted.
3x-2y=-18
4x+5y=22
multiply the top by 5 and the bottom by 2
15x-10y=-90
8x+10y=44
add them
23x=-46
x=-2
y=6